OpenStudy (anonymous):

A projectile is fired in such a way that its horizontal range is equal to 12.5 times its maximum height. What is the angle of projection?

OpenStudy (akashdeepdeb):

Use the formulas!! That's it!! :D

OpenStudy (anonymous):

Im trying to figure this out. I'm using pretty common notation. Vyf, Vxf, Vyi, Vxi, R, and h

OpenStudy (anonymous):

I'm trying to use the formulas but idk how to connect the dots

OpenStudy (amistre64):

h = -gt^2/2 + Vo sin(a) t d = t cos(a)

OpenStudy (amistre64):

forgot a Vo in my d

OpenStudy (amistre64):

the vertex of the parabola would be half the horizontal ... might want to keep that in the back of the picture

OpenStudy (akashdeepdeb):

R = u^2 sin 2theta/g H = u^2 sin^2 theta/2g R = 12.5 * H Now find theta!! :D Understood? :)

OpenStudy (akashdeepdeb):

http://en.wikipedia.org/wiki/Projectile_motion

OpenStudy (amistre64):

\[h=-\frac12gt^2+V_osin(\alpha)~t\] \[d=V_ocos(\alpha)~t=12.5h\]when:\[h=\frac{V_o}{12.5}cos(\alpha)~t\]

OpenStudy (amistre64):

\[h=-\frac12gt^2+V_osin(\alpha)~t\] \[h'=-gt+V_ocos(\alpha)=0~:~t=\frac{V_ocos(\alpha)}{g}\] \[h_{max}=-\frac12g\left(\frac{V_ocos(\alpha)}{g}\right)^2+V_osin(\alpha)~\frac{V_ocos(\alpha)}{g}\] \[d_{max}=V_ocos(\alpha)~t,~when~h=0\] \[0=t\left(-\frac12gt+V_osin(\alpha)\right)\] \[t=0, ~t=\frac{2~V_osin(\alpha)}{g}\] \[d_{max}=V_ocos(\alpha)~\frac{2~V_osin(\alpha)}{g}=12.5~h_{max}\]

OpenStudy (amistre64):

\[V_ocos(\alpha)~\frac{2~V_osin(\alpha)}{g}=-6.25g\left(\frac{V_ocos(\alpha)}{g}\right)^2+12.5V_osin(\alpha)~\frac{V_ocos(\alpha)}{g}\] \[V_ocos(\alpha)~\frac{2~V_osin(\alpha)}{g}-12.5V_osin(\alpha)~\frac{V_ocos(\alpha)}{g}=-6.25g\left(\frac{V_ocos(\alpha)}{g}\right)^2\] \[-\frac{10.5}{g}(V_o)sin(\alpha)cos(\alpha)=-6.25g\left(\frac{V_ocos(\alpha)}{g}\right)^2\] \[-{10.5}(V_o)sin(\alpha)cos(\alpha)=-6.25g^2\left(\frac{V_ocos(\alpha)}{g}\right)^2\] \[{10.5}(V_o)sin(\alpha)cos(\alpha)=6.25(V_o)^2 cos^2(\alpha)\] \[{10.5}sin(\alpha)cos(\alpha)=6.25V_o~cos^2(\alpha)\] \[{10.5}sin(\alpha)=6.25V_o~cos(\alpha)\]

OpenStudy (amistre64):

the Vo goes away, i missed it on the left above ....

OpenStudy (amistre64):

\[{10.5}sin(\alpha)=6.25~cos(\alpha)\] \[\frac{10.5}6.25tan(\alpha)=1\] etc..

OpenStudy (amistre64):

the latex went wonky on me .. (10.5/6.25) tan(a)=1 tan(a)= 6.25/10.5 a= arctan(6.25/10.5) maybe